OpenStudy (anonymous):

simplify [4 5] [1 8] [0 -1]+2[-4 -6]

OpenStudy (anonymous):

Multiply all elements in the right matrix by 2. Then add the matrix element wise. So the element in the upper left corner is going to 4+2*1=6. Can you find the other elements?

OpenStudy (anonymous):



OpenStudy (anonymous):

thats what i got 2 16 -8 -12

OpenStudy (anonymous):

right, now you need to add that to \[\left[\begin{matrix}4 & 5 \\ 0 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

add those by 2?

OpenStudy (anonymous):

you've multiplied the rigth matrix by 2, now you need to add the left matrix with matrix you just found, so: \[\left[\begin{matrix}4 & 5 \\ 0 & 1\end{matrix}\right]+\left[\begin{matrix}2 &16 \\ -8 &-12\end{matrix}\right]\]

OpenStudy (anonymous):

6 21 -8 -13 ?

OpenStudy (anonymous):

that's right. I see that I put 1 instead -1 in my matrices. Sorry about that.